OpenStudy (anonymous):

Find the equation of the line containing the given pair of points... (-7,-1) and (-3,-2)

OpenStudy (anonymous):

y = mx + b is line with m being slope and b is y-intercept. Slope of two points is delta of y / delta of x. So, here slope m = -1/10. y = -x/10 + b. Plugging in point -3,-2 into that equation, you get value of b. -2 = 3/10 + b => b = -23/10 So, equation is: y = -x/10 - 23/10

OpenStudy (anonymous):

thanks